• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name D.Wang/A.Camsonne

    Log entry time 03:50:54 on October19,2009

    Entry number 295608

    keyword=compton test runs summary

    moved Edet to -3.0 mm
    beam current 100uA
    FADC triggered on electron
    2 plane coincidence, or width 2.

    run20503 coincidence run, prescales 4, 0, 100. no coincidence events.

    run20504 coincidence run, prescales 0,0,100. still much fewer coincidence events than expected got recorded. when laser if off, there're coincidence events. When laser turns on, the DAQ seems to get stuck and no coincidence events got recorded.

    Alex suggested to change the "Polar gate" in spyconf from 27ms to 20 ms. Then the magic happened...

    run20505 coincidence run. prescale 0,0,100. Polar gate 20ms. short run. A lot of coincidence events got recorded. The "clock:numpolar" plot of h7 looks good. Correlation between photon ADC and electron strips is seen.

    run20506 coincidence run. prescales 1, 5, 100. Polar gate 20ms. longer run. Over 1M coincidence events. Coincidence is working at high current!

    So changing the parameter "Polar gate" in spyconf gets the coincidence events back for high current.

    run20507 photon only run with 15 mins laser off and 15 mins laser on.

    Polar gate set back to 27ms
    Edet moved to garage
    FADC trigger set back to standard
    Production run resumed.